4. Solve the following system of equations.

50π‘₯+20𝑦=1800 10π‘₯+3𝑦=300


1
Expert's answer
2022-03-24T16:08:34-0400

"\begin{cases}
 50x+20y=1800, \\
 10x+3y=300;
\end{cases} \implies"


"\begin{cases}
 50x+20y=1800, \\
 50x+15y=1500;
\end{cases} \implies"


"\begin{cases}
 50x+20y=1800, \\
 5y=300;
\end{cases}\implies"


"\begin{cases}
 50x+20y=1800, \\
 y=60;
\end{cases}\implies"


"\begin{cases}
 50x+1200=1800, \\
 y=60;
\end{cases} \implies"


"\begin{cases}
 x=12, \\
 y=60.
\end{cases}"